A 0.130-kg cube of ice (frozen water) is floating in glycerine. The gylcerine is in a tall cylinder that has inside radius 3.50 cm . The level of the glycerine is well below the top of the cylinder.

a) If the ice completely melts, by what distance does the height of liquid in the cylinder change? Express your answer with the appropriate units.

b) Does the level of liquid rise or fall? That is, is the surface of the water above or below the original level of the gylcerine before the ice melted?
Answer #1

Solution,

Density = m/V

Mass = density x volume

0.13 = 1.26 x v

Volume, V = 0.103 L

Difference in volume = 0.13 - 0.103 = 0.027

Volume, V = pi x r^2 h

Height, h = V/pi r^2 = 0.027/(pi x 0.35^2) = 0.070 = 0.70 cm higher

B)

Level will rise

Comment in case any doubt please rate my answer ...